# breaklines

### v3

@@ Abstract.
We leveraged AlphaTrans to translate ten real-world open-source projects consisting of $\langle$836836836836, $8575$, $2719$$\rangle$ (application and test) classes, (application and test) methods, and unit tests. AlphaTrans breaks down these projects into $17874$ fragments and translates the entire repository. $96.40\%$ of the translated fragments are syntactically correct, and AlphaTrans validates the translations’ runtime behavior and functional correctness for $27.03\%$ and $25.14\%$ of the application method fragments. On average, the integrated translation and validation take $34$ hours (min=$3$, max=$121$) to translate a project, showing its scalability in practice. For the syntactically or semantically incorrect translations, AlphaTrans generates a report including existing translation, stack trace, test errors, or assertion failures. We provided these artifacts to two developers to fix the translation bugs in four projects. They fixed the issues in $20.1$ hours on average ($5.5$ hours for the smallest and $34$ hours for the largest project) and achieved all passing tests. Without AlphaTrans, translating and validating such big projects could take weeks, if not months.
@@ 1. Introduction
This paper presents AlphaTrans, a neuro-symbolic [^1] The keyword symbolic here refers to a general term of symbolic learning in contrast to machine learning and should not be confused with symbolic execution. We refer to combining LLMs and program analysis as a neuro-symbolic approach. approach for automated repository-level code translation and validation. AlphaTrans leverages static analysis to resolve PL-specific features of the source language (§4.1), decompose the source project into smaller fragments (§4.2), and create a compilable project skeleton in the target language (§5). It then starts translating fragments in the reverse call order and validates them using existing tests when possible (§6). After translating each fragment, AlphaTrans updates the project skeleton and ensures the whole project compiles, gradually translating and validating the source project into the target PL. To improve translation quality, static analysis again comes to the rescue: AlphaTrans collects relevant context for each fragment, including translated callee methods and surrounding contexts, e.g., class declaration, global variables/fields, etc. It also uses relevant in-context examples based on the specific properties of the fragment to be translated. The current version of AlphaTrans implements two levels of dynamic validation: (1) running the source tests on the translated fragment in isolation using language interoperability (§6.1) and (2) decomposing, translating, and executing the unit tests on the translated project (§6.3). Finally, AlphaTrans recomposes the translated fragments to create the program in the target PL (§6.2).
@@
The idea of compositional translation and validation proposed by us is PL-agnostic; however, implementing the program transformation component is PL-specific. For the first version of AlphaTrans, the implementation supports translating from Java to Python. Our motivations for choosing this PL pair are: (1) Java offers many features that are not supported or common in other PLs by default (e.g., method/constructor overloading, complex types, circular dependencies, local or anonymous inner classes, interfaces, etc.); (2) Python programs are not compiled but interpreted, which makes many translation issues that can be caught at the compile time stay there until test execution and challenge the validation; and (3) both PLs are popular (top-5 on the TIOBE index (TIOBE, 2023)).
@@
Using AlphaTrans to translate ten real-world Java projects to Python corroborates its effectiveness: It can translate $17874$ field/method/test fragments, with $96.40\%$ syntactical correctness. For the $4654$ application method fragments that can be further evaluated through test execution, AlphaTrans achieves $27.03\%$ runtime validation and $25.14\%$ functional equivalence. AlphaTrans is scalable, completing translations in $34$ hours, on average. Human subjects improved partial translation of AlphaTrans and achieved green test suites within $20.1$ hours, on average, showing practicality of AlphaTrans. These results were achieved using a moderate-size open-access LLM (DeepSeek-Coder-33b-Instruct (Guo et al., 2024)). A stronger model, i.e., GPT-4o, improves the performance of AlphaTrans to $99.2\%$ syntactic correctness for all fragments and $27.95\%$ functional equivalence for application method fragments, with an overhead of \$$\$14.39$ per project, on average. The affordable cost is due to the novel features of the pipeline, namely, decomposition into fragments, prompt crafting, in-isolation validation of translations, and efficient feedback loop.
@@
To the best of our knowledge, AlphaTrans is the first technique to translate an entire repository, including tests, and generates validated translations (considering existing tests). The only prior repository-level translation attempt using GPT-4 (Pan et al., 2024) (translating Apache Commons CLI from Java to Python) resulted in non-compilable code, let alone the translation being validated. AlphaTrans is also the first technique leveraging language interoperability for in-isolation validation of translated fragments. The effort of human subjects to fix translation bugs by AlphaTrans and to achieve green tests creates pragmatic bug data sets for testing, bug localization, and program repair research. Our code and artifacts are publicly available for reproducing the results or translating new projects (Lab, 2025).
@@ 2. Challenges in Repository-Level Code Translation
Challenge 1: Class Size. The class consists of $25,380$ tokens ( a). Instructions for translating the code, in-context examples, and the model’s response can also significantly increase the number of input tokens. While some commercial LLMs support tens of thousands of tokens, many open-access LLMs do not. For example, DeepSeek-Coder-33b-Instruct (Guo et al., 2024) used in this paper has a context window of $16,384$ tokens, of which only $4,096$ tokens can be used for generation. To address this challenge, AlphaTrans decomposes Java application classes into smaller field or method fragments and translates each separately in a reverse call order (§4.2.1, §4.2.2).
@@
Challenge 2: PL-specific Properties. Java programs frequently use method/constructor overloading, which is not supported by default in Python ( a). This example shows instances of constructor overloading (lines 2 and 5). In Python, declaring two constructors is allowed; however, at runtime, the last declaration overrides all previous constructors, resulting in unexpected behavior. To address this issue, AlphaTrans employs program analysis to refactor the original code while preserving the functionality (through test execution). The transformation includes changing the constructor’s name, updating the references, and changing the constructor’s implementation. The transformed code ( b) makes the source program amenable to translation to Python.
@@
Challenge 3: Validation. To illustrate the challenges with validation, consider test01 ( c) that invokes four methods in its body (ArgExp, getNumArgs, getLen, and createMsg) to test the functionality of method getMsg in the assert statement. Suppose we can successfully translate all methods except createMsg. If we choose test translation, the most natural way of validating code translation, the execution of the translated test results in a runtime error when invoking createMsg. As a result, a translation issue in one method casts a shadow in validating the translation of the other methods. We refer to this issue as the test translation coupling effect. To overcome this challenge, AlphaTrans executes source language tests as-is (i.e., without translation) by leveraging a language-interoperability framework called GraalVM (Oracle, 2024) ( f). In this setting, a test in the source language is executed every time one of its invoked application methods (method fragments) is translated. This approach validates functional equivalence of each method in isolation since other methods invoked in the test or the body of the translated method can remain in the source language.
@@
Challenge 4: Test Translation. GraalVM has certain limitations (§6.1), which prevents AlphaTrans from validating all the code fragments in isolation. Furthermore, we need to translate tests regardless of whether they are used for validation to maintain the translated projects in the target language. Test errors as a result of test translation coupling effect under-approximates the quality of translation: failing to validate the translation of four methods because of one incorrect translation. Root causing the translation bugs also requires additional efforts from developers, i.e., looking at the stack trace and coverage. To overcome this challenge, AlphaTrans decomposes the original test suite into test fragments ( d). Executing the translated decomposed test suite results in three test passes ( e), validating the runtime behavior of three methods that the original test suite could not promptly provide.
@@
An alternative approach is parsing the stack trace and code coverage results for each runtime error during translation. However, test decomposition is a cleaner way to see the results per test execution promptly. It is also done once before translation. In translation to interpreted languages such as Python, specifically, the execution of test fragments can validate the runtime behavior of methods before waiting for functional validation. For fragments that GraalVM cannot validate, if AlphaTrans can successfully translate all the methods invoked during test execution and test passes, such test will also be used for validating functional correctness.
@@ 3. Overview of Approach
The type translation and skeleton construction phase takes the schema as input and produces target project skeleton, i.e., a compilable project in the target language with method signatures but no method implementation (§5.2). The first translation step happens here, where it translates the source PL types to the target PL to ensure that class skeletons are compilable (§5.1). The outcome of type translation is a type mapping from the source to the target PL, which AlphaTrans can reuse in translating other projects.
The compositional translation and validation takes schema and project skeleton as inputs and translates fragments, in the reverse call order, by prompting an LLM. After translating a fragment, it updates the class skeleton with a new translation and checks whether the skeleton is compilable. For method fragments, AlphaTrans looks for corresponding tests and, if any exist, validates them. The first level of validation is performed through GraalVM’s language interoperability to isolate the validation of the method using tests in the source language. Next, AlphaTrans translates and executes the corresponding tests. In case of compilation errors or test failures, AlphaTrans re-prompts the LLM with feedback (from the compilation and runtime errors) to improve the translation. If no improvement is achieved within a certain budget, AlphaTrans continues to the next fragment until all are translated. For methods whose translations are not compilable or result in test errors/failures, AlphaTrans generates reports consisting of existing translations and relevant artifacts, such as stack traces, test errors/failures, and test coverage information.
@@ 4.1. Program Transformation
This component performs semantics-preserving refactoring of method and constructor overloading in Java code to make it amenable to translation to Python. Other Java-specific features, namely, circular dependencies, inner classes, interfaces, and abstract classes, will be handled when constructing the project skeleton in Python (§5.2). The reason for resolving method and constructor overloading in the source language is that we have to change the implementation, i.e., call sites to methods and constructors. Therefore, such changes should be validated using source tests.
For overloaded methods, AlphaTrans makes each method name unique by adding an integer suffix (starting at 00) to the name and updates all call sites based on the new method names. Resolving overloaded constructors is not as straightforward, as they should have the same name as the enclosed declaring class. Furthermore, the invocation of constructors inside each other and the Java inheritance mechanism makes constructor overloading complex. Our algorithm (Algorithm 1) for resolving the constructor overloading handles three prominent [^2] Following the best practices for constructor overloading from Stack Overflow and analyzing the use of constructor overloading in open-source projects, we categorized the use cases into the mentioned rules. use cases shown in Figure 3.
@@
The first pattern (Figure 3-a) shows multiple independent constructors. AlphaTrans merges these constructors into one and uses an id parameter to differentiate between them. All the call sites of the constructors will be updated accordingly, based on the given id. The second use case (Figure 3-b) is more challenging to resolve, as one constructor calls the other using this(). AlphaTrans transforms the first constructor into a factory method and invokes the second constructor inside it. Factory methods are static, and AlphaTrans updates the call sites by directly invoking them on the class instance, e.g., ArgExp.ArgExp1(id,opt,msg). The last use case (Figure 3-c) is similar to the second, except that both constructors implement some code. AlphaTrans refactors the first constructor into a factory method and adds an extra id parameter to differentiate between behaviors implemented by different constructors. The call sites of the constructors will be updated like in previous cases. Real-world projects often combine these patterns, which AlphaTrans handles using Algorithm 1.
@@ 4.2.1. Source Decomposition
Real-world projects can include hundreds of files with thousands of lines of code, which exceed the context window of state-of-the-art LLMs. AlphaTrans employs static analysis to decompose code into smaller fragments, i.e., field fragments and method fragments. A field fragment includes modifiers, type, name, and field value. A field fragment can belong to an application or test class. A method fragment includes the full signature of an individual method in the source language. A method fragment can be an application or test method (e.g., helper methods or unit tests). During decomposition, AlphaTrans extracts meta-information related to the fragments, such as their location in code (e.g., start and end line), code (e.g., implementation between start and end line), dependencies (e.g., callers and callees), types (of inputs and output), and other necessary information such as file paths, class inheritance, imports, and method annotations. AlphaTrans stores fragments and their corresponding collected meta-data in a data structure called schema, which components in other phases will use. AlphaTrans also extracts the call graph to guide the translation, i.e., translate fragments in reverse call order.
@@ 4.2.2. Test Decomposition
The burden of validating functional equivalence in AlphaTrans is on GraalVM. Yet, we still need to translate and execute tests to validate the fragments that GraalVM cannot be used to validate (§6.1). Unit tests in real-world projects can invoke multiple methods and include multiple assert statements. Furthermore, long call chains are inevitable in real-world projects due to the high degree of intra- and iter-procedural dependencies. As we show later (§7.4), the average number of direct method invocations and method executions in tests for our studies subjects are $3$ and $27$, respectively. This can result in test translation coupling effect discussed in §2.
To enable test translation for validating runtime validation or functional equivalence, AlphaTrans decomposes each unit test into a series of test fragments, as shown in Figure 1-d. It uses each statement with a call to an application method as a cutting point. For statements enclosed by branches, loops, of exception-handling blocks, AlphaTrans includes the entire block. This process generates an ordering of executable test fragments for each unit test. Each test fragment includes all the statements of the lower-order fragments, along with additional statements that invoke only one method that was not invoked by previous fragments. AlphaTrans executes test fragments in increasing order until a test fails and skips running following fragments, as they will also fail.

### v5

# AlphaTrans: A Neuro-Symbolic Compositional Approach for Repository-Level Code Translation and Validation

@@ Abstract.
We leveraged AlphaTrans to translate ten real-world open-source projects consisting of $\langle$836836836836, $8575$, $2719$$\rangle$ (application and test) classes, (application and test) methods, and unit tests. AlphaTrans breaks down these projects into $17874$ fragments and translates the entire repository. $96.40\%$ of the translated fragments are syntactically correct, and AlphaTrans validates the translations’ runtime behavior and functional correctness for $27.03\%$ and $25.14\%$ of the application method fragments. On average, integrated translation and validation takes $34$ hours (min=$3$, max=$121$) to translate a project, showing its scalability in practice. For the syntactically or semantically incorrect translations, AlphaTrans generates a report including existing translation, stack trace, test errors, or assertion failures. We provided these artifacts to two developers to fix the translation bugs in four projects. They fixed the issues in $20.1$ hours on average ($5.5$ hours for the smallest and $34$ hours for the largest project) and achieved all passing tests. Without AlphaTrans, translating and validating such big projects could take weeks, if not months.
@@
copyright: ccdoi: 10.1145/3729379journalyear: 2025journal: PACMSEjournalvolume: 2journalnumber: FSEarticle: FSE109publicationmonth: 7ccs: Software and its engineering Source code generationccs: Computing methodologies Machine translation
@@ 1. Introduction
This paper presents AlphaTrans, a neuro-symbolic [^1] The keyword symbolic here refers to a general term of symbolic learning in contrast to machine learning and should not be confused with symbolic execution. We refer to combining LLMs and program analysis as a neuro-symbolic approach. approach for automated repository-level code translation and validation. AlphaTrans leverages static analysis to resolve PL-specific features of the source language (§4.1), decompose the source project into smaller fragments (§4.2), and create a compilable project skeleton in the target language (§5). It then starts translating fragments in reverse call order and validates them using existing tests when possible (§6). After translating each fragment, AlphaTrans updates the project skeleton and ensures the whole project compiles, gradually translating and validating the source project into the target PL. To improve translation quality, static analysis again comes to the rescue: AlphaTrans collects relevant context for each fragment, including translated callee methods and surrounding contexts, e.g., class declaration, global variables/fields, etc. It also uses relevant in-context examples based on the specific properties of the fragment to be translated. AlphaTrans implements two types of dynamic validation: (1) running the source tests on the translated fragments in isolation using language interoperability (§6.1) and (2) decomposing, translating, and executing the source tests on the translated fragments (§6.3). Finally, AlphaTrans recomposes the translated fragments to create the program in the target PL (§6.2).
@@
Our approach of compositional translation and validation is PL-agnostic; however, implementing the program transformation component is PL-specific. For the first version of AlphaTrans, the implementation supports translating Java code to Python. Our motivations for choosing this PL pair are: (1) Java offers many features that are not supported or common in other PLs by default (e.g., method/constructor overloading, complex types, circular dependencies, local or anonymous inner classes, interfaces, etc.); (2) Python programs are not compiled but interpreted, which makes many translation issues that can be caught at the compile time remain undetected until test execution and challenge the validation; and (3) both PLs are popular (top-5 on the TIOBE index (TIOBE, 2024)).
@@
Using AlphaTrans to translate ten real-world Java projects to Python corroborates its effectiveness: It can translate $17874$ field/method/test fragments, with $96.40\%$ syntactic correctness. For the $4654$ application method fragments that can be further evaluated through test execution, AlphaTrans achieves $27.03\%$ runtime validation and $25.14\%$ functional equivalence using the source tests. AlphaTrans is scalable, completing translations in $34$ hours, on average. Human subjects improved partial translation of AlphaTrans and achieved passing test suites within $20.1$ hours, on average, showing practicality of AlphaTrans. These results were achieved using a moderate-size open-access LLM (DeepSeek-Coder-33b-Instruct (Guo et al., 2024)). A stronger model, i.e., GPT-4o, improves the performance of AlphaTrans to $99.2\%$ syntactic correctness for all fragments and $27.95\%$ functional equivalence for method fragments, with an overhead of \$$\$14.39$ per project, on average. The affordable cost is due to the novel features of the pipeline, namely, decomposition into fragments, prompt crafting, in-isolation validation of translations, and efficient feedback loop.
@@
To the best of our knowledge, AlphaTrans is the first technique to translate an entire repository, including tests, and generates validated translations (considering existing tests). The only prior repository-level translation attempt using GPT-4 (Pan et al., 2024) (translating Apache Commons CLI from Java to Python) resulted in non-compilable code, let alone the translation being validated. AlphaTrans is also the first technique leveraging language interoperability for in-isolation validation of translated fragments. The effort of human subjects to fix translation bugs by AlphaTrans and achieve passing tests creates pragmatic bug data sets for testing, bug localization, and program repair research. Our code and artifacts are publicly available for reproducing the results or translating new projects (Lab, 2025a).
@@ 2. Challenges in Repository-Level Code Translation
Challenge 1: Class Size. The class consists of $25,380$ tokens ( a). Instructions for translating the code, in-context examples, and the model’s response can also significantly increase the number of input tokens. While some commercial LLMs support tens of thousands of tokens, many open-access LLMs do not. For example, DeepSeek-Coder-33b-Instruct (Guo et al., 2024) used in this paper has a context window of $16,384$ tokens, of which only $4,096$ tokens can be used for generation. To address this challenge, AlphaTrans decomposes Java application classes into smaller field and method fragments and translates each separately in reverse call order (§4.2.1, §4.2.2).
@@
Challenge 2: PL-specific Properties. Java programs frequently use method/constructor overloading, which is not supported directly in Python ( a). This example shows instances of constructor overloading (lines 2 and 5). In Python, declaring two constructors is allowed; however, at runtime, the last declaration overrides all previous constructors, resulting in unexpected behavior. To address this issue, AlphaTrans employs program analysis to refactor the original code while preserving the functionality (through test execution). The transformation includes changing the constructor’s name, updating the references, and changing the constructor’s implementation. The transformed code ( b) makes the source program amenable to translation to Python.
@@
Challenge 3: Validation. To illustrate the challenges with validation, consider test01 ( c) that invokes four methods in its body (ArgExp, getNumArgs, getLen, and createMsg) to test the functionality of method getMsg in the assert statement. Suppose we can successfully translate all methods except createMsg. If we choose test translation (a natural way of validating code translation), the execution of the translated test results in a runtime error when invoking createMsg. As a result, a translation issue in one method casts a shadow in validating the translation of the other methods. We refer to this issue as the test translation coupling effect. To overcome this challenge, AlphaTrans executes source language tests as-is (i.e., without translation) by leveraging a language-interoperability framework called GraalVM (Oracle, 2024) ( f). In this setting, a test in the source language is executed each time one of its invoked application methods (method fragments) is translated. This approach validates functional equivalence of each method in isolation as the other invoked application methods during test execution remain in the source language.
@@
Challenge 4: Test Translation. GraalVM has certain limitations (§6.1), which prevents AlphaTrans from validating all the code fragments in isolation. Furthermore, we need to translate tests regardless of whether they are used for validation to maintain the translated projects in the target language. Test errors due to test translation coupling effect under-approximate the quality of translation: failing to validate the translation of four methods because of one incorrect translation. To overcome this challenge, AlphaTrans decomposes the original test suite into test fragments ( d). Executing the translated decomposed test suite results in three test passes ( e), validating the runtime behavior of three methods that the original test suite could not promptly provide.
@@
An alternative approach is parsing the stack trace and code coverage results for each runtime error during translation. However, test decomposition is a cleaner way to see the results per test execution promptly. It is also done once before translation. In translation to interpreted languages such as Python, specifically, the execution of test fragments can validate the runtime behavior of methods before waiting for functional validation. For fragments that GraalVM cannot validate, if AlphaTrans can successfully translate all the methods invoked during test execution and the test passes, such a test will also be used for validating functional correctness.
@@ 3. Overview of Approach
The type translation and skeleton construction phase takes the schema as input and produces target project skeleton, i.e., a compilable project in the target language with method signatures but no method implementation (§5.2). The first translation step happens here, where the source PL types are translated to the target PL to ensure that class skeletons are compilable (§5.1). The outcome of type translation is a type mapping from the source to the target PL, which AlphaTrans can reuse in translating other projects.
The compositional translation and validation phase takes the schema and project skeleton as inputs and translates fragments, in reverse call order, by prompting an LLM. After translating a fragment, it updates the class skeleton with a new translation and checks whether the skeleton compiles. For a method fragment, AlphaTrans looks for corresponding tests and, if any exist, uses them to validate the fragment. The first level of validation is performed through GraalVM’s language interoperability to isolate the validation of the method using tests in the source language. Next, AlphaTrans translates and executes the tests. In case of compilation errors or test failures, AlphaTrans reprompts the LLM with feedback (from the compilation/runtime errors) to improve the translation. If no improvement is achieved within a certain budget, AlphaTrans continues to the next fragment until all are translated. For methods whose translations are not compilable or result in test errors/failures, AlphaTrans generates reports consisting of existing translations and relevant artifacts, such as stack traces, test errors/failures, and test coverage information.
@@ 4.1. Program Transformation
This component performs semantics-preserving refactoring of method and constructor overloading in Java code to make it amenable to translation to Python. Other Java-specific features, namely, circular dependencies, inner classes, interfaces, and abstract classes, are handled while constructing the project skeleton in Python (§5.2). The reason for resolving method and constructor overloading in the source language is that we have to change the implementation, i.e., call sites to methods and constructors. Therefore, such changes should be validated using source tests before translation.
For overloaded methods, AlphaTrans makes each method name unique by adding an integer suffix (starting at 00) to the name, and updates all call sites based on the new method names. Resolving overloaded constructors is not as straightforward, as they should have the same name as the enclosed declaring class. Furthermore, the invocation of constructors inside each other and the Java inheritance mechanism makes constructor overloading complex. Our algorithm (Algorithm 1) for resolving the constructor overloading handles three prominent patterns shown in Figure 3.222Following the best practices for constructor overloading from Stack Overflow and analyzing the use of constructor overloading in open-source projects, we categorized the use cases into the three patterns.
@@
The first pattern (Figure 3-a) shows multiple independent constructors. AlphaTrans merges these constructors into one and uses an id parameter to differentiate between them. All call sites of the constructors are updated accordingly to use the appropriate id value. The second pattern (Figure 3-b) involves a constructor call chain using this(). AlphaTrans transforms the first constructor into a factory method and invokes the second constructor inside it. Factory methods are static, and AlphaTrans updates the call sites to invoke them directly on the class, e.g., ArgExp.ArgExp1(id,opt,msg). The last pattern (Figure 3-c) is similar to the second one, except that both constructors implement some code. AlphaTrans refactors the first constructor into a factory method and adds an extra id parameter to differentiate between behaviors implemented by different constructors. The constructor call sites are updated accordingly, as in the previous cases. Real-world projects often combine these patterns, which AlphaTrans handles using Algorithm 1.
@@ 4.2.1. Source Decomposition
Real-world projects can include hundreds of files with thousands of lines of code, which exceed the context window of state-of-the-art LLMs. AlphaTrans employs static analysis to decompose code into smaller fragments, i.e., field fragments and method fragments. A field fragment includes modifiers, type, name, and field value. A field fragment can belong to an application or test class. A method fragment includes the method signature and can be an application or test method (e.g., helper methods or unit tests). During decomposition, AlphaTrans extracts meta-information related to the fragments, such as their location (e.g., start and end line), code (e.g., implementation between start and end line), dependencies (e.g., callers and callees), types (of inputs and output), and other necessary information such as file paths, class inheritance, imports, and method annotations. AlphaTrans stores fragments and their corresponding collected meta-data in a data structure called schema, which is used in the other phases. AlphaTrans also extracts the call graph to guide the translation, i.e., to translate fragments in reverse call order.
@@ 4.2.2. Test Decomposition
The burden of checking functional equivalence in AlphaTrans is on GraalVM. Yet, we still need to translate and execute tests to validate the fragments that cannot be validated by GraalVM (§6.1). Unit tests in real-world projects can invoke multiple methods and include multiple assert statements. Furthermore, long call chains are inevitable in real-world projects due to the high degree of intra- and iter-procedural dependencies. As we show later (§7.4), the average number of direct method invocations and method executions in tests for our studies subjects are $3$ and $27$, respectively. This can result in test translation coupling effect, discussed in §2.
To enable test translation for runtime validation or checking functional equivalence, AlphaTrans decomposes each unit test into a series of test fragments, as shown in Figure 1-d. It uses each statement with a call to an application method as a cutting point. For statements enclosed by branches, loops, of exception-handling blocks, AlphaTrans includes the entire block. This process generates an ordering of executable test fragments for each unit test. Each test fragment includes all the statements of the lower-order fragments, along with additional statements that invoke one additional method not invoked by previous fragments. AlphaTrans executes test fragments in increasing order until a test fails and skips running following fragments, as they will also fail.